# Semi-empirical Modeling of Supermassive Black Hole Evolution: Highlighting a possible tension between Demographics and Gravitational Wave Background

### v1

@@ 2 Semi-empirical modeling
A very effective approach to study the statistical evolution of the supermassive BH population in a data-driven way relies on the continuity equation [119, 120, 121, 122, 123, 124, 23, 24, 25, 125]. This allows to self-consistently trace the cosmic evolution of the BH mass function and the associated fraction of active BHs, building on the amount of mass accreted at each cosmic epoch. The latter is extracted from the input AGN luminosity function, from the adopted relation between AGN luminosity and BH mass as specified by the Eddington ratio distribution of active BHs, and from the assumed radiative (plus kinetic) efficiency. Moreover, being based on a linear differential equation, this approach allows to transparently disentangle the relative roles of accretion and binary mergers as complementary growth channels of the BH population at each cosmic time.
@@
The first addendum on the right hand side is a ‘drift’ term describing growth by gas accretion via a mean accretion rate $\langle\dot{M}_{\bullet}\rangle$ averaged over the BH population; the underlying rationale is that, although individual BHs may turn on and off, the statistical evolution depends only on the mean accretion rate as a function of mass and time. The second addendum $[\partial_{t}N]_{\bullet\bullet\rightarrow\bullet}$ is a ‘coagulation’ term accounting for binary mergers between BHs; note that mergers do not add net mass to the BH population but can alter the mass function by redistribution of the masses, from the low to the high mass end [^1] Note, however, that numerically solving the continuity equation on a finite grid of BH masses requires to assume that a constant flux of sub-resolution masses filtrates through the lowest mass boundary, see [126].. We now turn to describe these two ingredients in some more detail.
2.1 Modeling the growth by accretion
@@
We model the population growth by accretion basing on the redshift-dependent AGN bolometric luminosity function $N(\log L_{\rm AGN},t)$ derived from X-ray, mid-IR and optical data by [74] (details are provided in Appendix A). The AGN luminosity can in turn be related to the BH mass as $L_{\rm AGN}=\lambda\,M_{\bullet}\,c^{2}/t_{\rm Edd}$ in terms of the Eddington ratio $\lambda\equiv L_{\rm AGN}/L_{\rm Edd}$ and of a related timescale $t_{\rm Edd}=c\,\sigma_{\rm T}/4\pi G\,\mu_{e}m_{p}\approx 0.45$ Gyr, dependent solely on fundamental constants [^2] The Eddington luminosity $L_{\rm Edd}\equiv M_{\bullet}c^{2}/t_{\rm Edd}=4\pi G\mu_{e}m_{p}\,c\,M_{\bullet}/\sigma_{\rm T}\approx 1.3\times 10^{38}\,(M_{\bullet}/M_{\odot})$ erg s-1 is the limiting value for which the continuum radiation force emitted by the accretion disk around the BH balances gravity in isotropic conditions; in the above definition, $G$ is the gravitational constant, $\mu_{e}$ is the mean molecular weight per electron, $m_{p}$ is the proton mass, $c$ is the speed of light, and $\sigma_{\rm T}$ is the Thomson cross section.. We allow for a distribution in Eddington ratio of the BH population; in particular, as suggested by observations in the local and distant Universe [64, 66, 67, 37, 68, 89, 90], we assume a lognormal shape [^3] We also try to adopt a Schechter-like or a lognormal$+$powerlaw shape, which have been sometime considered in the literature as inspired from theoretical arguments or empirical analysis of type-2 AGNs at relatively low $z$ (e.g., [22, 127, 123, 23, 24, 25]. This requires to add at least one parameter (e.g., powerlaw behavior at small BH masses and its redshift evolution), which however we find to be largely unconstrained given the datasets considered in this work. Thus we prefer to use the pure log-normal expression involving the minimal set of parameters. In any case, we checked that qualitatively our results concerning the buildup of the BH mass function, the role of BH mergers and their contribution to the GW background are unaffected by the choice of the Eddington ratio distribution shape; minor quantitative alterations pertain only to the evolution of the BH-accretion parameters.
@@
with dispersion $\sigma_{\log\lambda}\approx 0.35$ dex and with a redshift dependent mean
@@
The accretion rate associated to a given AGN luminosity can be expressed as $\dot{M}_{\bullet}=(1-\epsilon-\epsilon_{\rm kin})\,L_{\rm AGN}/\epsilon\,c^{2}$ where $\epsilon$ and $\epsilon_{\rm kin}$ are the radiative and kinetic (jets and/or winds) efficiency, respectively. For a standard thin disk the radiative efficiency may range from $\epsilon\sim 0.06$ for a non-rotating to $\epsilon\sim 0.42$ for a maximally spinning Kerr BH [128, 59]. However, both for high accretion rates with $\lambda\gtrsim$ a few and for very low accretion rates with $\lambda\lesssim 0.01$, the radiative efficiency tends to decrease due to the onset of an advection dominated regime. In the former case, mainly applying at high redshift, the inflowing gas has a huge density and provides a very large optical depth to trap the radiation and advect it back into the BH [129, 130, 131]. In the latter case, mainly applying to low redshift and in the re-activations of dormant BHs, the inflowing gas density is so low to hinder cooling within an accretion timescale, and to store the viscous energy as a thermal component that is eventually advected into the hole [132, 133, 134]. We have devised a handy $\lambda-$dependent expression for the radiative efficiency as a function of the Eddington ratio that describes these various regimes:
@@
with $\epsilon_{\rm thin}$ a fitting parameter representing the efficiency of standard disk accretion; the quantities $\lambda_{c}\approx 0.01$ and $s\approx 1$ determine the behavior in the low accretion regime as suggested by the aforementioned studies, and are held fixed. The above detailed expression is adopted for the sake of completeness, but we anticipate that the advection dominated regimes will turn out to be marginally relevant for the evolution of the BH mass function. This is because the Eddington ratio distribution parameters inferred from our analysis will imply that the thin disk regime is the most typical accretion mode, with mildly super-Eddington values limited to high redshift $z\gtrsim 4-6$ and considerably sub-Eddington ones applying only to low $z\lesssim 1$ and large BH masses. As for the kinetic efficiency due to winds and jets, we assume $\epsilon_{\rm kin}\approx 0.15\,\epsilon_{\rm thin}$ as suggested by [135]; we anticipate that given the average values $\epsilon_{\rm thin}\lesssim 0.1$ inferred from our analysis, this position implies rather limited kinetic efficiencies. However, note that a few literature studies [136, 137] claim rather higher $\epsilon_{\rm kin}\lesssim 0.2$, though not applying to all the BH population; in any case, we will explore such a possibility in Appendix B. Given the above, the only free parameter of our model specifying the relation between luminosity and accretion rate is the thin-disk radiative efficiency $\epsilon_{\rm thin}$.
The mean accretion rate $\langle\dot{M}_{\bullet}\rangle$ entering Equation (2.1) is given by
@@
where $N_{\rm AGN}(\log M_{\bullet},t)\equiv\langle\delta\rangle(M_{\bullet},t)\,N(\log M_{\bullet},t)$ is the active BH mass function and $\langle\delta\rangle(M_{\bullet},t)$ is the mean BH duty cycle. The rationale of this expression is to weight the accretion rate of a BH by the duty cycle and to average it over the Eddington ratio distribution. The active BH mass function is in turn linked to the observed AGN luminosity function by the convolution
this equation may be solved parametrically by assuming a standard double power-law shape [^4] Although assuming a double power-law shape for the active BH mass function is standard practice, it is worth making a disclaimer here. There is no question that the AGN luminosity function is well described by a double power-law behavior. Since we have assumed a lognormal shape for the Eddington ratio distribution to reflect the latest observational determinations, we are naturally led to adopt a double power-law shape for the active mass function, otherwise the convolution Equation (2.5) would not produce the correct shape to fit the luminosity functions. On the other hand, we point out that such an assumption has the direct consequence of yielding a power-law shape in the relic BH mass function at the high-mass end. Despite this is somewhat supported by the data (see Section 3.1), the presence of an exponential falloff is still not excluded. Forcing the accretion model to produce it could be achieved by employing a Schechter fit for the active mass function. However, this solution would reduce the high-mass end of the predicted BH mass function at all redshifts, so lowering the BH merger rates and exacerbating the tension with the GW background (see Section 3.1). [138, 124, 23] for the active mass function $N_{\rm AGN}=N_{0}/[(M_{\bullet}/M_{\bullet,0})^{\omega_{1}}+(M_{\bullet}/M_{\bullet,0})^{\omega_{2}}]$ and then by fitting for the best parameters $(N_{0},M_{\bullet,0},\omega_{1},\omega_{2})(t)$ at any redshift to reproduce the AGN luminosity function. Note that another way to proceed will be to physically model the duty cycle and the Eddington ratio distribution by assuming an average light-curve shape for active BHs, as it has been pursued in [22, 123, 24]; however, in the spirit of a data-driven model here we do not take such a route.
@@
which recovers the simplest solution pioneered by [121, 122].
2.2 Modeling the growth by mergers
The most general framework [^5] In the literature, the evolution of the BH mass function has been approached with a Fokker-Planck (diffusion-like) equation [139, 140, 23] but this is only an approximation. In fact, it can be shown (for details, see [141]) that the Smoluchowski equation reduces to a Fokker-Planck equation when the mass distribution is sharply peaked around a mean value and the system is in a regime where small changes in mass dominate. These conditions, however, are not guaranteed to hold during the overall evolution of the BH population. Moreover, the Fokker-Planck approximation is not appropriate to distinguish between the production and destruction terms due to binary mergers; an unambiguous assessment of the former is instead required to estimate the GW background. to describe the evolution of the mass distribution $N(M_{\bullet},t)$ for a population of BHs under binary aggregations is provided by the Smoluchowski equation [142, 143, 144, 145, 146, 108]:

### v2

@@ 2.1 The continuity equation
As introduced above, a very effective approach to study the statistical evolution of the supermassive BH population in a data-driven way relies on the continuity equation [Cavaliere1971, Small1992, Marconi2004, Shankar2004, Yu2008, Cao2010, Shankar2013, Aversa2015, Tucci2017, Sicilia2022]. This method allows to self-consistently trace the cosmic evolution of the BH mass function and the associated fraction of active BHs, building on the amount of mass accreted at each cosmic epoch. The latter is extracted from the input AGN luminosity function, from the adopted relation between AGN luminosity and BH mass as specified by the Eddington ratio distribution of active BHs, and from the assumed radiative (plus kinetic) efficiency. Moreover, being based on a linear differential equation, this approach allows to transparently disentangle the relative roles of accretion and binary mergers as complementary growth channels of the BH population at each cosmic time.
@@
The first addendum on the right hand side is a ‘drift’ term describing growth by gas accretion via a mean accretion rate $\langle\dot{M}_{\bullet}\rangle$ averaged over the BH population; the underlying rationale is that, although individual BHs may turn on and off, the statistical evolution depends only on the mean accretion rate as a function of mass and time. The second addendum $\partial_{t}N]_{\bullet\bullet\rightarrow\bullet}$ is a ‘coagulation’ term accounting for binary mergers between BHs; note that mergers do not add net mass to the BH population but can alter the mass function by redistribution of the masses, from the low to the high mass end [^1] Note, however, that numerically solving the continuity equation on a finite grid of BH masses requires to assume that a constant flux of sub-resolution masses filtrates through the lowest mass boundary, see [Shankar2009].. We now turn to describe these two ingredients in some more detail.
2.2 Growth by accretion
To compute the growth by accretion we rely on one crucial observational input, constituted by the redshift-dependent AGN bolometric luminosity function $N(\log L_{\rm AGN},t)$ derived from multi-wavelength data in optical, UV, IR and X-ray bands by [Shen2020]. Specifically, we use the analytic rendition by [Shen2020] labeled ’A’ in their Table 4. This is expressed as
$N(\log L_{\rm AGN})=\frac{\phi_{\star}}{(L_{\rm AGN}/L_{\star})^{\gamma_{1}}+(L_{\rm AGN}/L_{\star})^{\gamma_{2}}}$ (2.2)
in terms of the parameters $\{\phi_{\star}(z),\gamma_{1}(z),\gamma_{2}(z),L_{\star}(z)\}$. The latter depend on redshift $z$ according to the expressions
$\log\phi_{\star}\,[{\rm Mpc}^{-3}\,{\rm dex}^{-1}]=d_{0}+d_{1}\,(1+z)$ (2.3)
$\gamma_{1}=a_{0}+a_{1}\,(1+z)+a_{2}\,[2\,(1+z)^{2}-1]$
$\gamma_{2}=\frac{2\,b_{0}}{[(1+z)/3]^{b_{1}}+[(1+z)/3]^{b_{2}}}$
$\log L_{\star}\,[L_{\odot}]=\frac{2\,c_{0}}{[(1+z)/3]^{c_{1}}+[(1+z)/3]^{c_{2}}}$
where $\{a_{0},a_{1},a_{2}\}=\{0.8569,-0.2614,0.02\}$, $\{b_{0},b_{1},b_{2}\}=\{2.5375,-1.0425,1.1201\}$, $\{c_{0},c_{1},c_{2}\}=\{13.0088,-0.5759,0.4554\}$, $\{d_{0},d_{1}\}=\{-3.5426,-0.3936\}$. The resulting bolometric AGN luminosity functions are plotted as a function of redshift over the range $z\sim 0-7$ (color-coded) in Figure 2.
We exploit the AGN luminosity function to compute the accretion rate $\langle\dot{M}_{\bullet}\rangle$ averaged over the population of active BH, which enters in the drift term of the continuity Equation (2.1). To make the computation explicit, we need some preliminary definitions and steps. To start with, it is convenient to relate the AGN luminosity $L_{\rm AGN}$ to the BH mass $M_{\bullet}$. One defines the Eddington ratio $\lambda\equiv L_{\rm AGN}/L_{\rm Edd}$, which is the ratio between the actual AGN luminosity and the Eddington luminosity. The Eddington luminosity $L_{\rm Edd}\equiv 4\pi G\mu_{e}m_{p}\,c\,M_{\bullet}/\sigma_{\rm T}\approx 1.3\times 10^{38}\,(M_{\bullet}/M_{\odot})$ erg s-1 is the limiting value for which the continuum radiation force emitted by the accretion disk around the BH balances gravity in isotropic conditions; in compact form one can write $L_{\rm Edd}=M_{\bullet}c^{2}/t_{\rm Edd}$ in terms of the Eddington timescale $t_{\rm Edd}=c\,\sigma_{\rm T}/4\pi G\,\mu_{e}m_{p}\approx 0.45$ Gyr, only dependent on fundamental constants [^2] In this definition, $G$ is the gravitational constant, $\mu_{e}$ is the mean molecular weight per electron, $m_{p}$ is the proton mass, $c$ is the speed of light, and $\sigma_{\rm T}$ is the Thomson cross section.. Provided these definitions, $L_{\rm AGN}=\lambda\,M_{\bullet}\,c^{2}/t_{\rm Edd}$ is just a rewriting that by itself does not require any assumption and does not add any relevant information unless one specifies how the Eddington ratio (or the AGN luminosity to BH mass ratio) varies during the lifetime of an individual AGN. However, this rewriting is helpful in view of a statistical description like the one subtended by the continuity equation, because as shown below the demographics (i.e., the mass function) of active BHs can be fully characterized by the observed AGN luminosity function and by the redshift-dependent distribution of Eddington ratios.
@@ 2.2 Growth by accretion
In fact, as suggested by observations in the local and distant Universe [Kollmeier2006, Willott2010, Kelly2013, Farina2022, Wu2022, He2024, Lai2024, Burke2025, Zou2025], active BHs features a distribution of Eddington ratios with a lognormal shape [^3] We also try to adopt a Schechter-like or a lognormal++powerlaw shape, which have been sometime considered in the literature as inspired from theoretical arguments or empirical analysis of type-2 AGNs at relatively low $z$ (e.g., [Hopkins2006, Kauffmann2009, Yu2008, Shankar2013, Aversa2015, Tucci2017]. This requires to add at least one parameter (e.g., powerlaw behavior at small BH masses and its redshift evolution), which however we find to be largely unconstrained given the datasets considered in this work. Thus we prefer to use the pure log-normal expression involving the minimal set of parameters. In any case, we checked that qualitatively our results concerning the buildup of the BH mass function, the role of BH mergers and their contribution to the GW background are unaffected by the choice of the Eddington ratio distribution shape; minor quantitative alterations pertain only to the evolution of the BH-accretion parameters.
@@
with dispersion $\sigma_{\log\lambda}\approx 0.35$ dex and with a redshift dependent mean. For the latter we adopt the scaling
@@
Then the active BH mass function is linked to the observed AGN luminosity function and to the Eddington ratio distribution by the convolution
this equation may be solved parametrically by assuming a standard double power-law shape [^4] Although assuming a double power-law shape for the active BH mass function is standard practice, it is worth making a disclaimer here. There is no question that the AGN luminosity function is well described by a double power-law behavior. Since we have assumed a lognormal shape for the Eddington ratio distribution to reflect the latest observational determinations, we are naturally led to adopt a double power-law shape for the active mass function, otherwise the convolution Equation (2.8) would not produce the correct shape to fit the luminosity functions. On the other hand, we point out that such an assumption has the direct consequence of yielding a power-law shape in the relic BH mass function at the high-mass end. Despite this is somewhat supported by the data (see Section 3.1), the presence of an exponential falloff is still not excluded. Forcing the accretion model to produce it could be achieved by employing a Schechter fit for the active mass function. However, this solution would reduce the high-mass end of the predicted BH mass function at all redshifts, so lowering the BH merger rates and exacerbating the tension with the GW background (see Section 3.1). [Schulze2010, Cao2010, Shankar2013] for the active mass function $N_{\rm AGN}=N_{0}/[(M_{\bullet}/M_{\bullet,0})^{\omega_{1}}+(M_{\bullet}/M_{\bullet,0})^{\omega_{2}}]$ and then by fitting for the best parameters $(N_{0},M_{\bullet,0},\omega_{1},\omega_{2})(t)$ at any redshift to reproduce the AGN luminosity function. Note that another way to proceed will be to physically model the Eddington ratio distribution by assuming an average light-curve shape for active BHs, as it has been pursued in [Hopkins2006, Yu2008, Aversa2015]; however, in the spirit of a data-driven model here we do not take such a route.
To proceed further, we need the BH accretion rate associated to a given AGN luminosity. This is provided by $\dot{M}_{\bullet}=(1-\epsilon-\epsilon_{\rm kin})\,L_{\rm AGN}/\epsilon\,c^{2}$ where $\epsilon$ and $\epsilon_{\rm kin}$ are the radiative and kinetic (jets and/or winds) efficiency, respectively [^5] The radiative efficiency $\epsilon$ is conventionally defined with respect to the large-scale mass inflow rate $\dot{M}_{\rm inflow}$ so that $L_{\rm AGN}=\epsilon\,\dot{M}_{\rm inflow}c^{2}$, but the actual BH growth rate $\dot{M}_{\bullet}=(1-\epsilon-\epsilon_{\rm kin})\,\dot{M}_{\rm infl}$ is smaller because of radiative and kinetic losses. Putting together these expressions yields $\dot{M}_{\bullet}=(1-\epsilon-\epsilon_{\rm kin})\,L_{\rm AGN}/\epsilon\,c^{2}$.. Using the previous rewriting of $L_{\rm AGN}$ in terms of BH mass $M_{\bullet}$, this can also be expressed as $\dot{M}_{\bullet}=M_{\bullet}/\tau_{\rm ef}$ in terms of the $e$-folding time $\tau_{\rm ef}\equiv\epsilon\,t_{\rm Edd}/\lambda\,(1-\epsilon-\epsilon_{\rm kin})$, which would be the theoretical timescale required for the BH to exponentially grow $M_{\bullet}\propto e^{\tau/\tau_{\rm ef}}$ by a factor $e\approx 2.7$ if $\lambda$ were constant in time. For an average efficiency $\epsilon\sim 0.1$ and $\lambda\sim 1$ it takes on values $\tau_{\rm ef}\approx 5\times 10^{7}$ yr, which is often quoted in textbooks as the typical growth timescale for a BH. However, we stress that such a timescale does not directly correspond to the duty cycle or visibility timescale of an AGN, first because for individual AGNs $\lambda$ is not constant along the light-curve, second because the emission could be stochastic or strongly intermittent, and finally because the visibility time depends on environmental conditions (e.g., amount of obscuration), on the observational band, and on selection effects (see [AlonsoTetilla2025]). For example, literature estimates of the visibility time amount to $\lesssim$ a few $10^{7}$ yr in the optical band for unobscured quasars, while they can be as high as $\gtrsim 10^{8}$ yr in the hard X-ray band for heavily obscured AGNs [Shankar2004, Lapi2006].
@@
Some discussion on the radiative efficiency $\epsilon$ is in order here. For a standard thin disk the radiative efficiency may range from $\epsilon\sim 0.06$ for a non-rotating to $\epsilon\sim 0.42$ for a maximally spinning Kerr BH [Thorne1974, Davis2011]. However, both for high accretion rates with $\lambda\gtrsim$ a few and for very low accretion rates with $\lambda\lesssim 0.01$, the radiative efficiency tends to decrease due to the onset of an advection dominated regime. In the former case, mainly applying at high redshift, the inflowing gas has a huge density and provides a very large optical depth to trap the radiation and advect it back into the BH [Abramowicz1988, Watarai2000, Madau2014]. In the latter case, mainly applying to low redshift and in the re-activations of dormant BHs, the inflowing gas density is so low to hinder cooling within an accretion timescale, and to store the viscous energy as a thermal component that is eventually advected into the hole [Narayan1995, Blandford2004, Merloni2008]. We have devised a handy $\lambda$-dependent expression for the radiative efficiency as a function of the Eddington ratio that describes these various regimes:
@@
with $\epsilon_{\rm thin}$ a fitting parameter representing the efficiency of standard disk accretion; the quantities $\lambda_{c}\approx 0.01$ and $s\approx 1$ determine the behavior in the low accretion regime as suggested by the aforementioned studies, and are held fixed. The above detailed expression is adopted for the sake of completeness, but we anticipate that the advection dominated regimes will turn out to be marginally relevant for the evolution of the BH mass function. This is because the Eddington ratio distribution parameters inferred from our analysis will imply that the thin disk regime is the most typical accretion mode, with mildly super-Eddington values limited to high redshift $z\gtrsim 4-6$ and considerably sub-Eddington ones applying only to low $z\lesssim 1$ and large BH masses. As for the kinetic efficiency due to winds and jets, we assume $\epsilon_{\rm kin}\approx 0.15\,\epsilon_{\rm thin}$ as suggested by [Shankar2008]; we anticipate that given the average values $\epsilon_{\rm thin}\lesssim 0.1$ inferred from our analysis, this position implies rather limited kinetic efficiencies. However, note that a few literature studies [Ghisellini2013, Zubovas2018] claim rather higher $\epsilon_{\rm kin}\lesssim 0.2$, though not applying to all the BH population; in any case, we will explore such a possibility in Appendix B. Given the above, the only free parameter of our model specifying the relation between luminosity and accretion rate is the thin-disk radiative efficiency $\epsilon_{\rm thin}$.
We have now all the ingredients to compute the mean accretion rate $\langle\dot{M}_{\bullet}\rangle$ entering Equation (2.1). This is given by
@@
where $p(\log\lambda)$ is the Eddington ratio distribution from Equation (2.4), $N_{\rm AGN}(\log M_{\bullet},t)$ is the active BH mass function from Equation (2.6) and $\epsilon$ is the $\lambda$-dependent radiative efficiency from Equation (2.7). To better catch the meaning of this expression, note that the quantity $\lambda\,(1-\epsilon-\epsilon_{\rm kin})\,M_{\bullet}/\epsilon\,t_{\rm Edd}=M_{\bullet}/\tau_{\rm ef}=\dot{M}_{\bullet}$ is just the accretion rate $\dot{M}_{\bullet}$ of an individual BH, and the ratio $N_{\rm AGN}(\log M_{\bullet},t)/N(\log M_{\bullet},t)\equiv\langle\delta\rangle(M_{\bullet},t)$ between the mass function of active and overall BH population defines the so called mean duty cycle $\langle\delta\rangle(M_{\bullet},t)$. Thus the above expression can be also written in compact form as $\langle\dot{M}_{\bullet}\rangle=\int{\rm d}\log\lambda\;p(\log\lambda)\dot{M}_{\bullet}\langle\delta\rangle$, highlighting that the underlying rationale is just to weight the accretion rate of a BH by the duty cycle and to average it over the Eddington ratio distribution.
Note that in the continuity equation only the combination $\langle\dot{M}_{\bullet}\rangle N(\log M_{\bullet})=\int{\rm d}\log\lambda\,$ $p(\log\lambda)\,\dot{M}_{\bullet}\,N_{\rm AGN}$ enters the drift term describing accretion, and this depends just on the Eddington ratio distribution and the active BH mass function, not on the duty cycle. In fact, in our semi-empirical framework the duty cycle $\langle\delta\rangle(\log M_{\bullet},z)$ is not modeled ab-initio and does not serve as an input (apart possibly for the initial condition, see Section 2.4), but it is a derived outcome that can be computed after solving the continuity equation for the total BH mass function. However, a detailed comparison of the duty cycle so obtained with observations is not straightforward. First, because observational estimates of the duty cycle as a function of BH mass are appreciably uncertain, especially toward intermediate to high redshift. Second, because the estimates of the duty cycle depend on the observational band and on selection effects, like the luminosity threshold for an AGN to be considered as active. Nevertheless, we anticipate that in Section 3.1 our semi-empirical framework will be shown to produce local mean duty cycles in qualitative agreement with current observations.
@@
which recovers the simplest solution pioneered by [Marconi2004, Shankar2004].
2.3 Growth by mergers
The most general framework [^6] In the literature, the evolution of the BH mass function has been approached with a Fokker-Planck (diffusion-like) equation [Yu2002, Vittorini2005, Shankar2013] but this is only an approximation. In fact, it can be shown (for details, see [Leyvraz2003]) that the Smoluchowski equation reduces to a Fokker-Planck equation when the mass distribution is sharply peaked around a mean value and the system is in a regime where small changes in mass dominate. These conditions, however, are not guaranteed to hold during the overall evolution of the BH population. Moreover, the Fokker-Planck approximation is not appropriate to distinguish between the production and destruction terms due to binary mergers; an unambiguous assessment of the former is instead required to estimate the GW background. to describe the evolution of the mass distribution $N(M_{\bullet},t)$ for a population of BHs under binary aggregations is provided by the Smoluchowski coagulation equation [Smoluchowski1916, Cavaliere1992, Benson2005, Erickcek2006, Ellis2024, SatoPolito2025]: